Identify the unknown isotope X in each of the following decays. Enter X in the form Y-A, where Y is the one or two letter chemical symbol and A is the mass number. (a) x 224 Ra a (b) x 232- y Th (c) x 12 e- v (d) 51cr e x V
Answer #1

1) 228Th

In this reaction X's mass number = 224 + 4 = 228

atomic number = 88 + 2

= 90

2) 232Th

In this reaction mass and atomic number does not change.


3) 12B

In this reaction X's mass number = 12

atomic number = 6 - 1

= 5


4) 51Mn

In this reaction X's mass number = 51

atomic number = 24 + 1

= 25